If f(x) = e^(2x) what is f(1/4)? I do not understand how to find this.
Answer by scott8148(5880)

you just substitute the given value of x into the function ... f(1/4)=e^(2*(1/4)) ... f(1/4)=e^(1/2)

f(1/4)=sqrt(e) ... f(1/4) approximately equals 1.65